48dimes more than quarters. Value od dimes is equal to the value of quarters. How many dimes are there?
Answer by stanbon(75887)

48dimes more than quarters. Value of dimes is equal to the value of quarters. How many dimes are there?
----
Equations:
Quantity:: d = q + 48
Value:: 10d = 25q
--------
Substitute for "d" and solve for "q"::
10(q+48) = 25q
10q + 480 = 25q
15q = 480
-----
Ans: q = 32 (# of quarters)
d = q + 48 = 32+48 = 80 (# of dimes)
--------------
Cheers,
Stan H.
------------